A train is moving on a track at 30m/s a ball is thrown from it perpendicular to the direction of motion with 30m/s at 45°from horizontal find the distance of ball from the point of projection on train to the point where it strikes

Grade:11

4 Answers

Eshan
5 years ago
Dear student,

Time of flight of the motion=\dfrac{2usin\theta}{g}=3\sqrt{2}s

Net speed of ball in the horizontal plane=\sqrt{u_x^2+u_z^2}=\sqrt{(30cos45^{\circ})^2+30^2}=15\sqrt{6}m/s
Hence the total distance travelled=15\sqrt{6}m/s\times 3\sqrt{2}s=30\sqrt{3}m
Arihant Debnath
4 years ago
Range  = (u^2sin2theta)/g
u = 30m/s
theta = 45 degrees with the horizontal which is perpendicular to the motion on the train.
Thus, the distance of ball from the point of projection on the ground when it strikes the ground is 90 m.
karan
3 years ago
the ans is 90m bcoz range of ball will always be some with frame of train if we find from ground frame then the range will change to 90 root 3